OpenStudy (anonymous):

i need help figuring out how to solve -10y^3+25y^2+30y^4=0...been stuck on this problem for a half hour plz help

myininaya (myininaya):

5y^2(2y+5+6y^2)=0 5y^2=0 implies y=0 so find the solutions to 6y^2+2y+5=0 and you are done

OpenStudy (anonymous):

thats where im stuck at

OpenStudy (anonymous):

i do not understand this at all a whole test of problems like this

myininaya (myininaya):

use quadratic formula or you can use completing the square

OpenStudy (anonymous):

hmm confused lol suck at math

myininaya (myininaya):

\[y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

our equation is in the form ax^2+bx+c=0 what is a? what is b? what is c?

myininaya (myininaya):

and just pretend those x's are y's

OpenStudy (anonymous):

o ic

OpenStudy (anonymous):

ty trying to figure it all out

myininaya (myininaya):

a=6 b=2 c=5 so we have \[y=\frac{-2 \pm \sqrt{2^2-4(6)(5)}}{2(6)}=\frac{-2 \pm \sqrt{4-120}}{12}=\frac{-2 \pm \sqrt{-116}}{12}\] \[=\frac{-2 \pm \sqrt{-1*4*29}}{12}=\frac{-2 \pm \sqrt{-1}\sqrt{4}\sqrt{29}}{12}=\frac{-2 \pm i*2*\sqrt{29}}{12}\] \[=\frac{-2}{12} \pm i \frac{2\sqrt{29}}{12}=\frac{-1}{6} \pm i \frac{\sqrt{29}}{6}\]

OpenStudy (anonymous):

tyvm

myininaya (myininaya):

np